-3x-y=3 -3x-5y=-21
Answer by lynnlo(4176)

is this 1 problem or 2
i solved it as 1 problem
-3x-y=3
-3x-5y=21
answer y=-9/2,x=1/2
if this is 2 different problems send back and i will re-do